HUBBLE SPACE TELESCOPE DAILY REPORT #2194 PERIOD COVERED: 0000Z (UTC) 08/24/98 - 0000Z (UTC) 08/25/98 Daily Status Report as of 237/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. As described in 2.1 and HSTAR 6635, the acquisition for the fourth iteration of this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of NIC/1/2/3 7850 (Cosmological Parameters Omega and Lambda from High-Z Type Ia Supernovae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the type Ia super novae SN988 and SN98142 in order to obtain the cosmological parameters Omega and Lambda. HST can directly measure the cosmological parameters, Omega, Lambda, and thus the curvature, Omega_k, using Type Ia supernovae {SNe Ia} as calibrated standard candles. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Seven Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As described in 2.1 and HSTAR 6635, the acquisition for the sixth iteration of this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were completed as planned, and no further problems were reported. 1.5 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Seven Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.7 Completed NIC/2 7879 (Nuclear Rings in the IR: Hidden Super Star Clusters) The Near Infrared Camera (Camera Two) was used to make observations of the galaxy NGC-5248 in order to search in the infrared for star bursting nuclear rings. The rings contain large populations of super star clusters similar to those recently discovered in other types of starburst systems. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Nine Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.9 Completed STIS/CCD 7672 (STIS CCD Sensitivity Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star AGK+81D266 in order to monitor the sensitivity of each CCD grating mode to detect any change due to contamination or other causes. As described in 2.1 and HSTAR 6635, the acquisition for this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were completed as scheduled, and no further anomalies were noted. 1.10 Completed Three Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. As described in 2.1 and HSTAR 6635, the acquisition for the first iteration of this proposal defaulted to fine lock on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.11 Completed STIS/CCD/MA1/MA2 7733 (Distant Futures of Solar Activity) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the K2 III class star HD124897 (a.k.a., Arcturus). This is to explore possible future fates of solar magnetic activity through high-S/N ultraviolet spectra of this ancient Sun analog. The observations were completed as planned, and no anomalies were reported. 1.12 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.13 Completed NIC/1/2/3 7869 (The Morphology of Dense Gas in Seyferts , Obstruction and Fueling of AGNs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Seyfert galaxy NGC-5506 to study the morphology of its dense gas. Near-Infrared color maps and images in H-2 emission will be used to trace hot and cold dense gas in this Seyfert galaxy. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 As documented in HSTAR 6635, the acquisition as 236/144538Z defaulted to fine lock on FG-3 only when FGS-1 exceeded the scan step limit. The proposals described in 1.1, 1.4, 1.9, and 1.10 may have been affected. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: The SSR EDAC error counter was cleared four times using ROP SR-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Another iteration of the IP command test will be conducted this afternoon. Continuation of normal science observations and calibrations. /DMH